# On the microlocal phase space concentration of Schrödinger evolutions

### v1

@@ Abstract.
In this work, we investigate the microlocal properties of the evolutions of Schrödinger equations using metaplectic Wigner distributions. So far, only restricted classes of metaplectic Wigner distributions, satisfying particular structural properties, have allowed the analysis of microlocal properties. We first extend the microlocal results to all metaplectic Wigner distributions, including the well-known Kohn-Nirenberg quantization, and examine these findings in the framework of Fourier integral operators with quadratic phase. Finally, we analyze the phase space concentration of the (cross) Wigner distribution arising from the interaction of two states, with particular attention to interactions generated by certain Schrödinger evolutions. These contributions enable a more refined study of the so-called ghost frequencies.
@@ 1. Introduction
Microlocal analysis and time-frequency analysis have long been regarded as two complementary perspectives on phase space analysis, offering distinct yet interconnected approaches and tools to the study of operators and singularities in phase space. On the one hand, microlocal analysis is based on the notion of wave front set, introduced by Hörmander with the aim of analyzing the microlocal $\mathcal{C}^{\infty}$ and Sobolev singularities of distributions, and the directions in which these singularities occur, cf. [hormander1, Chapter 8] and [hormander3, Chapter 18]. Later, in [hormander1495quadratic], he introduced a different notion of wave front set, denoted by $WF_{\mathrm{iso}}$, tailored to detect Schwartz singularities. These singularities are called isotropic because the variables $x$ and $\xi$ are treated symmetrically. This concept has been extensively studied by many authors from both pseudodifferential and time-frequency analysis perspectives (see e.g. [cappiellohanarmonic, cappielloFios, MPT2025, PravdaRodinoWahlberg, wahlberg2018propagation, MR4182467]). In the latter context, the notion is commonly referred to as the Gabor wave front set and it is given via the short-time Fourier transform (STFT) [cordero2020time, grochenig2013foundations],
@@
Let $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ and $g\in\mathcal{S}(\mathbb{R}^{d})\setminus\{0\}$. We say that $0\neq X_{0}\notin WF_{G}(f)$, the Gabor wave front set of $f$, if there exists an open cone $\Gamma_{X_{0}}\subseteq\mathbb{R}^{2n}$ containing $X_{0}$ such that $V_{g}f\in L_{v_{s}}^{2}(\Gamma_{X_{0}})$ for every $s\geq 0$.
In [cordero2022wigner], a further alternative definition of the wave front set is proposed basically by replacing the STFT with the Wigner distribution [wigner1932quantum], here normalized as
@@
On the other hand, this qualitative approach of capturing singularities via wave front sets is reinterpreted by time-frequency analysis, where the STFT and the Wigner distribution are used as tools to quantify phase space concentration and how it changes through the action of evolution operators, pseudodifferential operators and, more generally, Fourier integral operators (FIOs), see [MR4182445, MR4286055, MR3571901, cordero2010time, MR3391896, MR3392649, MR2294795, MR2356420, MR2208883] and the references therein. In the Weyl-Wigner formulation of non-commutative quantum mechanics, the Wigner distribution is modified to encode the geometry induced by a canonical deformation of the Heisenberg algebra [MR2432025, MR2718929]. The Wigner distribution is a fundamental tool in both quantum mechanics and signal analysis, providing a powerful phase space representation of states and signals. Yet, in many concrete situations, ranging from quantum evolutions governed by non-standard symplectic forms to time-frequency transforms tailored to specific signal geometries, the classical Wigner distribution must be modified to properly capture the underlying structure [MR2843220, PrataDias2018]. Metaplectic Wigner distributions, recently defined by Cordero and Rodino, provide a fruitful framework encompassing all the main generalizations of the Wigner distribution in the literature. Their contributions also show that the Wigner distribution can effectively be used to study the evolution operators of Schrödinger equations, the so-called Wigner analysis. Ghost frequencies, however, arise naturally from the non-linear nature of the Wigner distribution. Alongside the development of Wigner analysis, the propagation of singularities measured by means of the Wigner distribution has been studied, even in a metaplectic framework. However, only restricted classes of metaplectic Wigner distributions, whose expressions can be derived explicitly, have been considered so far.
@@
We write $W_{\mathcal{A}}f:=W_{\mathcal{A}}(f,f)$. Many time–frequency representations that fall within the framework of metaplectic Wigner distributions have been extensively studied from various perspectives over the past decade. We refer the reader to the works [bayer2020linear, cordero2020linear, MR3471302, grochenig2024benedicks, grochenig2025characterization, grochenig2025more, Minh:2024aa, zhang2019uncertainty, zhang2021scaled], among the aforementioned works and others, though the literature is considerably broader. Our starting point is [cordero2024wigner], which ends with a brief discussion of $\mathcal{A}$-Wigner wave fronts, focusing on metaplectic Wigner distributions $W_{\mathcal{A}}$ satisfying certain structural properties, precisely covariance and shift-invertibility, we leave the details to Definitions 2.11 and 2.13 below, as they are secondary for the aim of this Introduction.
@@
(4) $\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}|W_{\mathcal{A}}f(X)|^{2}dX<\infty,$
where $\chi_{\Gamma_{X_{0}}}$ is the characteristic function on $\Gamma_{X_{0}}$.
@@
[cordero2024wigner, Theorem 7.6] Let $W_{\mathcal{A}}$ be a shift-invertible and covariant metaplectic Wigner distribution. Then, for every $a\in S^{0}_{0,0}(\mathbb{R}^{2n})$ and every $f\in L^{2}(\mathbb{R}^{n})$,
@@
The formal definition of $K$, investigated in [cordero2024understanding, cordero2023wigner], requires the use of the cross-Wigner distribution,
@@
see Definition 5.1 for the details. However, a finer analysis of these operators requires to restrict to the Wigner wave front set $WF_{\mathrm{w}}$ already considered in [cordero2022wigner].
@@
This notion provides a natural tool to describe correlations between states, how they mutually affect their singularities and eventually have insights on their propagation in phase space. Indeed, the interaction between two states produces the so-called ghost frequencies, fast oscillations that show up when computing the Wigner distribution of a signal, due to the absence of the mitigating effect of the window function present instead in Definition 1.1, see [cordero2024wigner, cordero2025wigner, cordero2022wigner] and the quadratic nature of the Wigner distribution. For the sake of clarity, consider a limit case where $f=1$ and $g=\delta$, the point-mass distribution. Then,
@@
The third addendum is the real part of $W(f,g)$ and produces oscillations that neither $f$ nor $g$ themselves exhibit, precisely the ghost frequencies [cordero2025wigner]. We are therefore naturally led to study the mutual interaction, from a microlocal perspective, of $u_{1}$ and $u_{2}$ Schrödinger evolutions of $u_{0,1},u_{0,2}\in L^{2}(\mathbb{R}^{n})$ determined by
@@
where, $t\in\mathbb{R}\mapsto S_{j,t}\in\operatorname*{Sp}(d,\mathbb{R})$ describes the solution of the classical equations of motion with Hamiltonian $a_{j}(x,\xi)$ in phase-space, and the symbols $b_{j,t}\in S^{0}_{0,0}(\mathbb{R}^{2n})$, see [cordero2014generalized]. In the case where $a_{1}=a_{2}$ the propagation of interferences between $u_{1}$ and $u_{2}$ is governed by the following result.
@@
We point out that theory stated in this work for operators having symbols in $S^{0}_{0,0}(\mathbb{R}^{2n})$ can be rephrased verbatim to far wider classes of symbols, the so-called weighted Sjöstrand classes [MR1266757]. The intersection of these spaces coincides with $S^{0}_{0,0}(\mathbb{R}^{2n})$, as it is detailed in [BastianoniDecay, Lemma 2.3], and eventually our results can be derived as particular instances of this more general framework. Our restriction to $S^{0}_{0,0}(\mathbb{R}^{2n})$ is therefore made purely for the sake of simplicity.
@@ 2.3. Metaplectic operators
Any such operator $\hat{S}$ is called a metaplectic operator. A careful analysis of (28) reveals that if it is satisfied by $\hat{S}$, then it is also satisfied by $c\hat{S}$ for every $c\in\mathbb{T}=\{z\in\mathbb{C}:|z|=1\}$ (phase factor). The group $\{\hat{S}:S\in\operatorname*{Sp}(n,\mathbb{R})\}$ has a subgroup consisting of exactly two metaplectic operators for each symplectic matrix, the so-called metaplectic group, denoted by $\operatorname*{Mp}(n,\mathbb{R})$. The projection $\pi^{Mp}:\hat{S}\in\operatorname*{Mp}(n,\mathbb{R})\mapsto S\in\operatorname*{Sp}(n,\mathbb{R})$ is a group homomorphism with kernel $\{\pm\operatorname*{id}_{L^{2}}\}$. For the theory of metaplectic operators we refer to [cordero2020time], but the interested reader may also refer to [de2021quantum, folland1989harmonic, grochenig2013foundations].
@@
Proposition 2.5 (Appendix B, [cordero2023symplectic]).
@@
Proposition 2.6 (Proposition A.2, [cordero2023metaplectic]).
@@
Finally, the following covariance property for the Wigner distribution (8) holds (see [lernermetrics], p. 61).
@@ 2.4. Metaplectic Wigner distributions
[cordero2024wigner, cordero2022wigner] Our theory is set in the framework of metaplectic Wigner distributions, a class of time-frequency representations generalizing the Wigner distribution using metaplectic operators. They were defined in [cordero2022wigner], and their properties were later established in [cordero2023symplectic, cordero2024excursus, cordero2024metaplectic, cordero2024wigner, cordero2024unified, cordero2023characterization, cordero2023metaplectic].
@@
    For $\tau=1/2$ and $\tau=0$ we retrieve the classical (cross-)Wigner distribution (8) and the Rihackzek distribution [Kirkhood1933, Rihaczek1968], respectively. Again, $W_{\tau}(f,g)=\widehat{\mathcal{A}_{\tau}}(f\otimes\bar{g})$, where
@@
    for a fixed $M\in\operatorname*{GL}(2d,\mathbb{R})$. If
@@
    see [cordero2024wigner], where these distributions are called totally Wigner decomposable.
@@
    The modified Wigner distributions in [PrataDias2018], arising when considering different symplectic forms on $\mathbb{R}^{2n}$, are rescaled Wigner distributions and, therefore, metaplectic Wigner distributions.
@@
    We also mention that many time-frequency representations of the metaplectic type were considered in the last decades, with applications to signal analysis, cf. [Bai2012, Minh1, Pei2001, ZhangLCWD, ZHANG2023108846, ZhangNewWigner] and the references therein.
@@
We also mention shift-invertible distributions, which play a fundamental role in classifying signals in terms of their time-frequency concentration, see [cordero2023symplectic] and [cordero2024metaplectic] for a more detailed discussion.
@@ 2.5. Pseudodifferential operators
[hormander3, lernermetrics] If $f,g\in\mathcal{S}(\mathbb{R}^{n})$, then $W(f,g)\in\mathcal{S}(\mathbb{R}^{2n})$. If $a\in\mathcal{S}^{\prime}(\mathbb{R}^{2n})$, we may consider the Weyl-quantized pseudodifferential operator (or, Weyl operator) with symbol $a$, that is the linear operator continuous from $\mathcal{S}(\mathbb{R}^{n})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ defined by
@@
We conclude this subsection by recalling two fundamental properties of Weyl operators: their metaplectic invariance and a composition result. We limit here to state these results in the form we need them, and we refer to [lernermetrics, Theorem 2.1.2] and [hormander3, Chapter 18], respectively, for a more detailed discussion.
@@ 2.6. Hamilton map
[hormandersymplectic] Let $a$ be a complex quadratic form on the phase space
@@ 3. $\mathcal{A}$-Wigner kernels
The microlocal analysis of pseudodifferential operators with metaplectic Wigner distributions requires the theory of $\mathcal{A}$-Wigner kernels, extending the already existing Wigner analysis considered in [cordero2024understanding, cordero2023wigner], limited to the Wigner distribution. Let us briefly synthesize the main results in this particular case.
@@ 3.1. Definition of $\mathcal{A}$-Wigner kernels and main properties
We have that $\mbox{span}(\{W_{\mathcal{A}}(f,g):f,g\in\mathcal{S}(\mathbb{R}^{n})\})$ is dense in $\mathcal{S}(\mathbb{R}^{2n})$, whereas $\mbox{span}(\{W_{\mathcal{A}}(u,v)\otimes\overline{W_{\mathcal{A}}(f,g)}:f,g,u,v\in\mathcal{S}(\mathbb{R}^{n})\})$ is dense in $\mathcal{S}(\mathbb{R}^{4n})$.
@@ 3.2. Doubling the variables
Stated differently, this problem is the same as defining an embedding $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})\hookrightarrow\mathcal{A}^{\prime}\in\operatorname*{Sp}(4d,\mathbb{R})$ properly so that the right side of (82) holds.
@@
(93) $\mathcal{D}_{P_{-}}(\mathcal{A}\otimes\bar{\mathcal{A}})$ $\mathcal{D}_{P_{+}}\!=\!\!\left(\begin{array}[]{cc|cc}UAU+R^{\top}AR&UAR^{\top}+R^{\top}AU^{\prime}&UBU-R^{\top}BR&UBR^{\top}-R^{\top}BU^{\prime}\\ RAU-U^{\prime}AR&RAR^{\top}-U^{\prime}AU^{\prime}&RBU+U^{\prime}BR&RBR^{\top}+U^{\prime}BU^{\prime}\\ \hline\cr UCU-R^{\top}CR&UCR^{\top}-R^{\top}CU^{\prime}&UDU+R^{\top}DR&UDR^{\top}+R^{\top}DU^{\prime}\\ RCU+U^{\prime}CR&RCR^{\top}+U^{\prime}CU^{\prime}&RDU-U^{\prime}DR&RDR^{\top}-U^{\prime}DU^{\prime}\end{array}\right)$

### v2

# On the microlocal phase-space concentration of Wigner distributions associated with Schrödinger evolutions

@@ Abstract.
In this work, we investigate the microlocal properties of the evolutions of Schrödinger equations using metaplectic Wigner distributions. So far, only restricted classes of metaplectic Wigner distributions, satisfying particular structural properties, have allowed the analysis of microlocal properties. We first extend the microlocal results to all metaplectic Wigner distributions, and examine these findings in the framework of Fourier integral operators with quadratic phase. Finally, we analyze the phase space concentration of the (cross) Wigner distribution arising from the interaction of two states, with particular attention to interactions generated by certain Schrödinger evolutions. These contributions enable a more refined study of the so-called ghost frequencies.
@@ 1. Introduction
Microlocal analysis and time-frequency analysis have long been regarded as two complementary perspectives on phase space analysis, offering distinct yet interconnected approaches and tools to the study of operators and singularities in phase space. On the one hand, microlocal analysis is based on the notion of wave front set, introduced by Hörmander with the aim of analyzing the microlocal $\mathcal{C}^{\infty}$ and Sobolev singularities of distributions, and the directions in which these singularities occur, cf. [46, Chapter 8] and [47, Chapter 18]. Later, in [44], he introduced a different notion of wave front set, denoted by $WF_{\mathrm{iso}}$, tailored to detect Schwartz singularities. These singularities are called isotropic because the variables $x$ and $\xi$ are treated symmetrically. This concept has been extensively studied by many authors from both pseudodifferential and time-frequency analysis perspectives (see e.g. [11, 12, 53, 56, 62, 63]). In the latter context, the notion is commonly referred to as the Gabor wave front set and it is given via the short-time Fourier transform (STFT) [28, 38],
@@
Let $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ and $g\in\mathcal{S}(\mathbb{R}^{n})\setminus\{0\}$. We say that $0\neq X_{0}\notin WF_{G}(f)$, the Gabor wave front set of $f$, if there exists an open cone $\Gamma_{X_{0}}\subseteq\mathbb{R}^{2n}$ containing $X_{0}$ such that $V_{g}f\in L_{v_{s}}^{2}(\Gamma_{X_{0}})$ for every $s\geq 0$.
In [29], a further alternative definition of the wave front set is proposed basically by replacing the STFT with the Wigner distribution [64], here normalized as
@@
On the other hand, this qualitative approach of capturing singularities via wave front sets is reinterpreted by time-frequency analysis, where the STFT and the Wigner distribution are used as tools to quantify phase space concentration and how it changes through the action of evolution operators, pseudodifferential operators and, more generally, Fourier integral operators (FIOs), see [6, 7, 13, 24, 25, 27, 39, 43, 60] and the references therein. In the Weyl-Wigner formulation of non-commutative quantum mechanics, the Wigner distribution is modified to encode the geometry induced by a canonical deformation of the Heisenberg algebra [3, 4]. The Wigner distribution is a fundamental tool in both quantum mechanics and signal analysis, providing a powerful phase space representation of states and signals. Yet, in many concrete situations, ranging from quantum evolutions governed by non-standard symplectic forms to time-frequency transforms tailored to specific signal geometries, the classical Wigner distribution must be modified to properly capture the underlying structure [35, 36]. Metaplectic Wigner distributions, recently defined by Cordero and Rodino, provide a fruitful framework encompassing all the main generalizations of the Wigner distribution in the literature. Their contributions also show that the Wigner distribution can effectively be used to study the evolution operators of Schrödinger equations, the so-called Wigner analysis. Ghost frequencies, however, arise naturally from the non-linear nature of the Wigner distribution. Alongside the development of Wigner analysis, the propagation of singularities measured by means of the Wigner distribution has been studied, even in a metaplectic framework. However, only restricted classes of metaplectic Wigner distributions, whose expressions can be derived explicitly, have been considered so far.
@@
We write $W_{\mathcal{A}}f:=W_{\mathcal{A}}(f,f)$. Many time–frequency representations that fall within the framework of metaplectic Wigner distributions have been extensively studied from various perspectives over the past decade. We refer the reader to the works [5, 31, 32, 40, 41, 42, 54, 66, 69], among the aforementioned works and others, though the literature is considerably broader. Our starting point is [20], which ends with a brief discussion of $\mathcal{A}$-Wigner wave fronts, focusing on metaplectic Wigner distributions $W_{\mathcal{A}}$ satisfying certain structural properties, precisely covariance and shift-invertibility, we leave the details to Definitions 2.11 and 2.13 below, as they are secondary for the aim of this Introduction.
@@
(4) $\int_{\Gamma_{X_{0}}}\langle X\rangle^{2s}|W_{\mathcal{A}}f(X)|^{2}dX<\infty.$
@@
[20, Theorem 7.6] Let $W_{\mathcal{A}}$ be a shift-invertible and covariant metaplectic Wigner distribution. Then, for every $a\in S^{0}_{0,0}(\mathbb{R}^{2n})$ and every $f\in L^{2}(\mathbb{R}^{n})$,
@@
The formal definition of $K$, investigated in [17, 19], requires the use of the cross-Wigner distribution,
@@
see Definition 5.1 for the details. However, a finer analysis of these operators requires to restrict to the Wigner wave front set $WF_{\mathrm{w}}$ already considered in [29].
@@
This notion provides a natural tool to describe correlations between states, how they mutually affect their singularities and eventually have insights on their propagation in phase space. Indeed, the interaction between two states produces the so-called ghost frequencies, fast oscillations that show up when computing the Wigner distribution of a signal, due to the absence of the mitigating effect of the window function present instead in Definition 1.1, see [20, 21, 29] and the quadratic nature of the Wigner distribution. For the sake of clarity, consider a limit case where $f=1$ and $g=\delta$, the point-mass distribution. Then,
@@
The third addendum is the real part of $W(f,g)$ and produces oscillations that neither $f$ nor $g$ themselves exhibit, precisely the ghost frequencies [21]. We are therefore naturally led to study the mutual interaction, from a microlocal perspective, of $u_{1}$ and $u_{2}$ Schrödinger evolutions of $u_{0,1},u_{0,2}\in L^{2}(\mathbb{R}^{n})$ determined by
@@
where, $t\in\mathbb{R}\mapsto S_{j,t}\in\operatorname*{Sp}(n,\mathbb{R})$ describes the solution of the classical equations of motion with Hamiltonian $a_{j}(x,\xi)$ in phase-space, and the symbols $b_{j,t}\in S^{0}_{0,0}(\mathbb{R}^{2n})$, see [23]. In the case where $a_{1}=a_{2}$ the propagation of interferences between $u_{1}$ and $u_{2}$ is governed by the following result.
@@
We point out that theory stated in this work for operators having symbols in $S^{0}_{0,0}(\mathbb{R}^{2n})$ can be rephrased verbatim to far wider classes of symbols, the so-called weighted Sjöstrand classes [59]. The intersection of these spaces coincides with $S^{0}_{0,0}(\mathbb{R}^{2n})$, as it is detailed in [2, Lemma 2.3], and eventually our results can be derived as particular instances of this more general framework. Our restriction to $S^{0}_{0,0}(\mathbb{R}^{2n})$ is therefore made purely for the sake of simplicity.
@@ 2.3. Metaplectic operators
Any such operator $\hat{S}$ is called a metaplectic operator. A careful analysis of (28) reveals that if it is satisfied by $\hat{S}$, then it is also satisfied by $c\hat{S}$ for every $c\in\mathbb{T}=\{z\in\mathbb{C}:|z|=1\}$ (phase factor). The group $\{\hat{S}:S\in\operatorname*{Sp}(n,\mathbb{R})\}$ has a subgroup consisting of exactly two metaplectic operators for each symplectic matrix, the so-called metaplectic group, denoted by $\operatorname*{Mp}(n,\mathbb{R})$. The projection $\pi^{Mp}:\hat{S}\in\operatorname*{Mp}(n,\mathbb{R})\mapsto S\in\operatorname*{Sp}(n,\mathbb{R})$ is a group homomorphism with kernel $\{\pm\operatorname*{id}_{L^{2}}\}$. For the theory of metaplectic operators we refer to [28], but the interested reader may also refer to [34, 37, 38].
@@
Proposition 2.5 (Appendix B, [14]).
@@
Proposition 2.6 (Proposition A.2, [58]).
@@
Finally, the following covariance property for the Wigner distribution (8) holds (see [52], p. 61).
@@ 2.4. Metaplectic Wigner distributions
[20, 29] Our theory is set in the framework of metaplectic Wigner distributions, a class of time-frequency representations generalizing the Wigner distribution using metaplectic operators. They were defined in [29], and their properties were later established in [14, 15, 16, 20, 22, 30, 58].
@@
    For $\tau=1/2$ and $\tau=0$ we retrieve the classical (cross-)Wigner distribution (8) and the Rihackzek distribution [50, 57], respectively. Again, $W_{\tau}(f,g)=\widehat{\mathcal{A}_{\tau}}(f\otimes\bar{g})$, where
@@
    for a fixed $M\in\operatorname*{GL}(2n,\mathbb{R})$. If
@@
    see [20], where these distributions are called totally Wigner decomposable.
@@
    The modified Wigner distributions in [36], arising when considering different symplectic forms on $\mathbb{R}^{2n}$, are rescaled Wigner distributions and, therefore, metaplectic Wigner distributions.
@@
    We also mention that many time-frequency representations of the metaplectic type were considered in the last decades, with applications to signal analysis, cf. [1, 51, 55, 65, 67, 68] and the references therein.
@@
We also mention shift-invertible distributions, which play a fundamental role in classifying signals in terms of their time-frequency concentration, see [14] and [16] for a more detailed discussion.
@@ 2.5. Pseudodifferential operators
[47, 52] If $f,g\in\mathcal{S}(\mathbb{R}^{n})$, then $W(f,g)\in\mathcal{S}(\mathbb{R}^{2n})$. If $a\in\mathcal{S}^{\prime}(\mathbb{R}^{2n})$, we may consider the Weyl-quantized pseudodifferential operator (or, Weyl operator) with symbol $a$, that is the linear operator continuous from $\mathcal{S}(\mathbb{R}^{n})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ defined by
@@
We conclude this subsection by recalling two fundamental properties of Weyl operators: their metaplectic invariance and a composition result. We limit here to state these results in the form we need them, and we refer to [52, Theorem 2.1.2] and [47, Chapter 18], respectively, for a more detailed discussion.
@@ 2.6. Hamilton map
[45] Let $a$ be a complex quadratic form on the phase space
@@ 3. $\mathcal{A}$-Wigner kernels
The microlocal analysis of pseudodifferential operators with metaplectic Wigner distributions requires the theory of $\mathcal{A}$-Wigner kernels, extending the already existing Wigner analysis considered in [17, 19], limited to the Wigner distribution. Let us briefly synthesize the main results in this particular case.
@@ 3.1. Definition of $\mathcal{A}$-Wigner kernels and main properties
The $\mbox{span}(\{W_{\mathcal{A}}(f,g):f,g\in\mathcal{S}(\mathbb{R}^{n})\})$ is dense in $\mathcal{S}(\mathbb{R}^{2n})$, whereas $\mbox{span}\{W_{\mathcal{A}}(u,v)\otimes\overline{W_{\mathcal{A}}(f,g)}:f,g,u,v\in\mathcal{S}(\mathbb{R}^{n})\}$ is dense in $\mathcal{S}(\mathbb{R}^{4n})$.
@@ 3.2. Doubling the variables
Stated differently, this problem is the same as defining an embedding $\mathcal{A}\in\operatorname*{Sp}(2n,\mathbb{R})\hookrightarrow\mathcal{A}^{\prime}\in\operatorname*{Sp}(4n,\mathbb{R})$ properly so that the right side of (82) holds.
@@
(89) $\mathcal{D}_{P_{-}}(\mathcal{A}\otimes\bar{\mathcal{A}})\mathcal{D}_{P_{+}}$
(94) $=\left(\begin{array}[]{cc|cc}UAU+R^{\top}AR&UAR^{\top}+R^{\top}AU^{\prime}&UBU-R^{\top}BR&UBR^{\top}-R^{\top}BU^{\prime}\\ RAU-U^{\prime}AR&RAR^{\top}-U^{\prime}AU^{\prime}&RBU+U^{\prime}BR&RBR^{\top}+U^{\prime}BU^{\prime}\\ \hline\cr UCU-R^{\top}CR&UCR^{\top}-R^{\top}CU^{\prime}&UDU+R^{\top}DR&UDR^{\top}+R^{\top}DU^{\prime}\\ RCU+U^{\prime}CR&RCR^{\top}+U^{\prime}CU^{\prime}&RDU-U^{\prime}DR&RDR^{\top}-U^{\prime}DU^{\prime}\end{array}\right)$